OpenStudy (samanthagreer):

Solve. A. s ≥ –2 B. s ≤ –2 C. s ≥ –18 D. s ≤ –18

OpenStudy (shinalcantara):

\[-\frac{ s }{ 3 } \ge 6\] cross multiplying 3 you'll have: \[-s \ge 3(6)\] \[-s \ge 18\] divide both sides with -1. doing so the inequality sign will also change \[s \le 18\]

OpenStudy (shinalcantara):

*-18

OpenStudy (shinalcantara):

yep

OpenStudy (shinalcantara):

\[\frac{ x }{ -2 } - 8 \le -2\] getting t\[\frac{ x -8(-2) }{ -2 } \le -2\] cross multiply -2 to the right side \[x + 16 \le 4\] \[x \le 4 -16\] \[x \le -12\]

OpenStudy (shinalcantara):

*nosebleeding

OpenStudy (shinalcantara):



OpenStudy (samanthagreer):

i figured it out